# A new framework for nonlinear Kalman filters

### v10

@@ Appendix C The detailed algorithm of various nonlinear Kalman filters used in the paper
32:     $\bm{\mathcal{X}}_{i,k|k}=\bm{\mathcal{X}}_{i,k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1}),i=0,\ldots,2n_{x}$
33:     $\bm{\mathcal{Y}}_{i,k|k}=\bm{h}(\bm{\mathcal{X}}_{i,k|k}),i=0,\ldots,2n_{x}$
34:     $\bm{\hat{y}}_{k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{Y}}_{i,k|k}$
35:     $\bm{P}_{xy,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k}-\bm{\hat{x}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
36:     $\bm{P}_{y,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
37:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
38:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$

### v11

@@ Appendix C The detailed algorithm of various nonlinear Kalman filters used in the paper
32:  $\bm{\mathcal{X}}_{i,k|k}=\bm{\mathcal{X}}_{i,k|k-1}+\bm{K}_{k}(\bm{z}_{k}-\bm{\hat{z}}_{k|k-1}),i=0,\ldots,2n_{x}$
33:  $\bm{\mathcal{Z}}_{i,k|k}=\bm{h}(\bm{\mathcal{X}}_{i,k|k}),i=0,\ldots,2n_{x}$
34:  $\bm{\hat{z}}_{k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{Z}}_{i,k|k}$
35:  $\bm{P}_{xz,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k}-\bm{\hat{x}}_{k|k}][\bm{\mathcal{Z}}_{i,k|k}-\bm{\hat{z}}_{k|k}]^{T}$
36:  $\bm{P}_{z,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{Z}}_{i,k|k}-\bm{\hat{z}}_{k|k}][\bm{\mathcal{Z}}_{i,k|k}-\bm{\hat{z}}_{k|k}]^{T}$
37:  $\bm{S}_{k|k}=\bm{P}_{z,k|k}+\bm{R}_{k}$
38:  $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}_{k}\bm{S}_{k|k}\bm{K}_{k}^{T}-\bm{P}_{xz,k|k}\bm{K}_{k}^{T}-\bm{K}_{k}\bm{P}_{xz,k|k}^{T}$
@@
Appendix D Supplementary results related to the “back out” step
Fig. 8, 9, 1517 compare the state estimation accuracy with and without the “back out” step. Additionally, Fig. 1822 quantify how often the proposed “back out” step is activated in practice by reporting the percentage of state updates that trigger this safeguard.
Figure 15: The root mean squared error of the state estimations with and without the “back out” step under different measurement noise setups (terrain-referenced navigation).
Figure 16: The root mean squared error of the state estimations with and without the “back out” step under different measurement noise setups (synchronous generator state estimation).
Figure 17: The root mean squared error of the state estimations with and without the “back out” step under different measurement noise setups (battery state estimation).
Figure 18: Percentage of state updates that trigger “back out” (3D target tracking).
Figure 19: Percentage of state updates that trigger “back out” (terrain-referenced navigation).
Figure 20: Percentage of state updates that trigger “back out” (synchronous generator state estimation).
Figure 21: Percentage of state updates that trigger “back out” (pendulum state estimation).
Figure 22: Percentage of state updates that trigger “back out” (battery state estimation).